# RefChartQA: Grounding Visual Answer on Chart Images through Instruction Tuning

### v1

@@ Abstract
Recently, Vision Language Models (VLMs) have increasingly emphasized document visual grounding to achieve better human-computer interaction, accessibility, and detailed understanding. However, its application to visualizations such as charts remains under-explored due to the inherent complexity of interleaved visual-numerical relationships in chart images. Existing chart understanding methods primarily focus on answering questions without explicitly identifying the visual elements that support their predictions. To bridge this gap, we introduce RefChartQA, a novel benchmark that integrates Chart Question Answering (ChartQA) with visual grounding, enabling models to refer elements at multiple granularities within chart images. Furthermore, we conduct a comprehensive evaluation by instruction-tuning 5 state-of-the-art VLMs across different categories. Our experiments demonstrate that incorporating spatial awareness via grounding improves response accuracy by over 15%, reducing hallucinations, and improving model reliability. Additionally, we identify key factors influencing text-spatial alignment, such as architectural improvements in TinyChart, which leverages a token-merging module for enhanced feature fusion. Our dataset is open-sourced for community development and further advancements. All models and code will be publicly available at https://github.com/moured/RefChartQA.
@@ 1 Introduction
Charts serve as an essential medium for conveying information in numerous documents, including scientific articles, business reports, and educational materials. Recent advances in chart understanding have led to promising progress in tasks such as chart visual question answering (Chart VQA) and chart-based information extraction [24, 14]. However, most existing works mainly concentrate on predicting the final answers based on charts, largely overlooking the interpretability aspect of chart understanding tasks. To address this, recent work on Visual Answer Grounding (VAG) extends VQA by highlighting the specific regions within an image that support the predicted answer[3, 34]. Although VAG has been explored in natural images, its application to data visualizations remains under-studied. In charts, grounding entails highlighting specific bars, points, or legend entries that justify a numerical or factual response [24].
@@
  • We introduce RefChartQA, a new dataset of real-world charts enriched with bounding-box annotations for answer grounding. It extends existing ChartQA resources [24] by focusing on questions that involve arithmetic or logical reasoning.
@@ 2.1 Chart QA Benchmarks
A variety of Chart QA benchmarks have been proposed to evaluate a model’s ability to analyze and reason over data visualizations. Initially, researchers relied on synthetic datasets like FigureQA [7] and DVQA [13], which use template-generated charts and questions. These studies demonstrated the feasibility of automating chart reasoning but often lacked the stylistic diversity and complexity of real-world charts. To address this, PlotQA [27] introduced open-vocabulary queries, capturing more advanced arithmetic or logical operations, yet remained inherently synthetic. Moving beyond templates, ChartQA [24] presented over 9,600 human-curated questions covering diverse chart types and domains. Subsequent endeavors have explored automated generation via LLMs, e.g., ChartX [30] and ChartLlama [8], or self-training methods, e.g., EvoChart [12] to broaden variety and complexity. Despite these steps, explicit visual grounding in chart contexts remains underexplored: only PlotQA includes limited bounding-box labels for basic tasks, highlighting a persistent need for benchmarks incorporating fine-grained answer localization.
@@ 2.2 Chart Understanding Models
Chart understanding has traditionally involved multi-stage pipelines: extracting textual or numerical content via OCR and component detectors, then converting that content into tabular form for downstream QA [27, 24]. While these methods yield interpretable intermediate results, the reliance on OCR can introduce errors, and the final QA often overlooks intricate visual details (colors, positions, etc.). More recently, end-to-end vision-language models have gained traction for chart tasks. Examples include UniChart [25], which adapts Donut [16] for chart comprehension, and Pix2Struct [17], which encodes visual features of chart screenshots for direct QA. State-of-the-art systems now harness instruction-tuned multimodal large language models like LLaVA [19], PaliGemma [1, 28], and ChartLlama [8] to process complex queries. Additionally, specialized techniques address numerical reasoning by coupling code execution and large language models, as demonstrated by TinyChart [32] and ChartGemma [26]. These approaches significantly boost answer accuracy but typically provide limited transparency about which regions of the chart were utilized for reasoning. Integrating explicit bounding-box-based visual grounding remains a gap, motivating our work to unify high-level chart QA performance with localized, interpretable evidence.
@@ 3 RefChartQA Benchmark
To enable visual grounding for chart images, we introduce RefChartQA, a benchmark that extends ChartQA [24] and TinyChart-PoT [32] with explicit grounding annotations. Unlike prior datasets that focus solely on answer accuracy, RefChartQA ensures models justify their predictions by linking answers to relevant chart elements. Grounding annotations are generated through a multi-stage process, including a heuristic-based single-element extraction, a Program-of-Thought (PoT)-based method, and a GPT-based method. The whole process is summarized in Fig. 3. Next, we discuss how we formed the annotations and present the statistical analysis of our benchmark.
Figure 3: Overview of the RefChartQA annotation process.
@@ 3.1 Annotation Process
We implement visual grounding using a bounding box paradigm, which is particularly effective for chart images due to its versatility across different chart types. However, we do not claim this to be the best or only approach; alternative methods, such as semantic segmentation or keypoint detection, could provide complementary advantages by capturing finer details or structural relationships within charts. These methods, however, often come with higher computational costs. For instance, autoregressive segmentation typically requires significantly more processing power and training time. While the keypoint paradigm is a more native approach, it is less suited for many common visualizations where answers correspond to regions rather than single points. Future work could explore multi-grounding paradigms that adapt based on the image and query context.
We choose ChartQA [24] as the foundation for our dataset due to its diverse question types, covering both arithmetic and non-arithmetic reasoning over chart images. Arithmetic questions often require multi-element operations, such as calculating averages or making comparisons, while non-arithmetic questions typically refer to a single chart element, though some may involve multiple elements for contextual understanding. ChartQA provides rich metadata, including bounding boxes for all text elements and their corresponding textual content. Additionally, we incorporate a subset of TinyChart’s ChartQA-PoT dataset [32], as it provides additional arithmetic questions based on ChartQA charts and structured reasoning in the form of Python code. However, both datasets lack explicit connections between visual elements and the questions, as they do not include direct grounding annotations. To bridge this gap, we develop a structured annotation process that links questions to their supporting visual evidence as summarized in Fig. 3.
@@
Single-Element Extraction. The first stage of the annotation pipeline targets single-element non-arithmetic questions, where the answer corresponds directly to a single data element in the chart. Since these questions do not involve computational reasoning or the interaction of multiple elements, a heuristic-based approach is used to match answers with their corresponding visual elements. For example, the question "What is the number of Macy’s stores worldwide in the year 2018?" describes a data retrieval operation which can be answered using structural information only. To identify the relevant evidence element, we utilize a string-matching technique that extracts x-axis labels and category names from the question and answer text. Since ChartQA questions frequently reference these elements explicitly, string matching provides a robust method for identifying the referenced part. The goal is to reduce the set of candidates to exactly one element, which would be the grounding element. If this fails, the question progresses to the next stage.
@@
The PoT process above starts by extracting the relevant metadata needed to answer the question—specifically, the list of years X and the number of people doing freelance work each year Y. This is followed by a sequence of reasoning operations, culminating in the final computed answer, which is stored in the variable Answer. To ensure consistency with ChartQA annotations, we first align the extracted list values with the chart’s metadata (JSON file). This step establishes a mapping between approximated values and metadata values while also verifying the accuracy of the underlying reasoning process. To further minimize annotation errors, we enforce that the PoT prediction must exactly match the ground truth answer. After validation, we extract the relevant reasoning elements using a template-based approach. The corresponding metadata elements serve as grounding annotations. If validation fails or the reasoning complexity exceeds the capabilities of the template-based approach, the question advances to the next stage.
@@ 5.1 Model Selection
To evaluate the impact of different pretrained models, we selected several SOTA MLLMs across chart analysis, document understanding, and visual grounding. For chart-related models, we chose UniChart [25], ChartGemma [26], and TinyChart [32]. UniChart, a lightweight Donut-based [16] model with 260M parameters, delivers competitive performance despite its smaller size. In contrast, ChartGemma and TinyChart ($\sim$3B parameters) incorporate advanced reasoning capabilities. TinyChart supports two inference modes: direct inference for end-to-end predictions and PoT-based reasoning, which generates executable Python code. ChartGemma, however, exclusively relies on PoT-based predictions. All three models have been instruction-tuned on chart-related tasks like summarization, chart-to-table conversion, and chart VQA. For visual grounding, we selected Qwen-VL-Chat [29], one of the few VG models fine-tuned on ChartQA. Alongside image captioning, VQA, and OCR, it has been trained on grounded image captioning and object referral, making it a strong candidate for grounding-based evaluations. For document understanding, we included DocOwl 1.5 [10], which, like chart-focused models, has been trained on document parsing, table extraction, chart parsing, and VQA, while also incorporating text localization capabilities.
5.2 Experiment Setup
We conducted our experiments on four Nvidia A40 GPUs with 48 GB of VRAM. All models were fine-tuned by parameter-efficient fine-tuning using LoRA adapters [11] for five epochs with a global batch size of 16. For UniChart, we used full-parameter instead of LoRA tuning due to its comparatively small parameter count. At this point, we treat the bounding box predictions as text and use the next-token prediction objective, with a softmax cross-entropy loss. For inference we relied on direct inference mode and the following input prompt (with slight variances depending on how the model processes images):
@@ 5.2 Experiment Setup
For UniChart, a special ground token task was introduced:
"<<<grounding_chartqa>>>{question} <<<s_answer>>>"
5.3 Evaluation Metrics
@@
[…]
@@ 6 Conclusion
In this paper, we introduced RefChartQA, the first benchmark that integrates chart question answering with visual answer grounding to enhance interpretability and response reliability in chart-based reasoning. Through comprehensive experiments on five state-of-the-art vision-language models across different categories, we demonstrate that incorporating grounding improves response accuracy, while also identifying key factors influencing text-spatial alignment. Despite these advancements, challenges remain in handling complex multi-element reasoning and numerical inference. We hope this work will drive further research in chart understanding and VAG, fostering more robust, interpretable, and spatially aware document understanding models.

### v2

@@ Abstract
Recently, Vision Language Models (VLMs) have increasingly emphasized document visual grounding to achieve better human-computer interaction, accessibility, and detailed understanding. However, its application to visualizations such as charts remains under-explored due to the inherent complexity of interleaved visual-numerical relationships in chart images. Existing chart understanding methods primarily focus on answering questions without explicitly identifying the visual elements that support their predictions. To bridge this gap, we introduce RefChartQA, a novel benchmark that integrates Chart Question Answering (ChartQA) with visual grounding, enabling models to refer elements at multiple granularities within chart images. Furthermore, we conduct a comprehensive evaluation by instruction-tuning 6 state-of-the-art VLMs across different categories. Our experiments demonstrate that incorporating spatial awareness via grounding improves response accuracy by over 15%, reducing hallucinations, and improving model reliability. Additionally, we identify key factors influencing text-spatial alignment, such as architectural improvements in TinyChart, which leverages a token-merging module for enhanced feature fusion. Our dataset is open-sourced for community development and further advancements. All models and code will be publicly available at https://github.com/moured/RefChartQA.
@@ 1 Introduction
Charts serve as an essential medium for conveying information in numerous documents, including scientific articles, business reports, and educational materials. Recent advances in chart understanding have led to promising progress in tasks such as chart visual question answering (Chart VQA) and chart-based information extraction [28, 15, 22]. However, most existing works mainly concentrate on predicting the final answers based on charts, largely overlooking the interpretability aspect of chart understanding tasks. To address this, recent work on Visual Answer Grounding (VAG) extends VQA by highlighting the specific regions within an image that support the predicted answer[4, 38]. Although VAG has been explored in natural images, its application to data visualizations remains under-studied. In charts, grounding entails highlighting specific bars, points, or legend entries that justify a numerical or factual response [28].
@@
  • We introduce RefChartQA, a new dataset of real-world charts enriched with bounding-box annotations for answer grounding. It extends existing ChartQA resources [28, 36] by focusing on questions involve arithmetic or logical reasoning.
@@ 2.1 Chart QA Benchmarks
A variety of Chart QA benchmarks have been proposed to evaluate a model’s ability to analyze and reason over data visualizations. Initially, researchers relied on synthetic datasets like FigureQA [8], DVQA [14] and [19], which use template-generated charts and questions. These studies demonstrated the feasibility of automating chart reasoning but often lacked the stylistic diversity and complexity of real-world charts. To address this, PlotQA [31] introduced open-vocabulary queries, capturing more advanced arithmetic or logical operations, yet remained inherently synthetic. Moving beyond templates, ChartQA [28] presented over 9,600 human-curated questions covering diverse chart types and domains. Subsequent endeavors have explored automated generation via LLMs, e.g., ChartX [34] and ChartLlama [9], or self-training methods, e.g., EvoChart [13] to broaden variety and complexity. Despite these steps, explicit visual grounding in chart contexts remains underexplored: only PlotQA includes limited bounding-box labels for basic tasks, highlighting a persistent need for benchmarks incorporating fine-grained answer localization.
@@ 2.2 Chart Understanding Models
Chart understanding has traditionally involved multi-stage pipelines: extracting textual or numerical content via OCR and component detectors, then converting that content into tabular form for downstream QA [31, 28, 21]. While these methods yield interpretable intermediate results, the reliance on OCR can introduce errors, and the final QA often overlooks intricate visual details (colors, positions, etc.). More recently, end-to-end vision-language models have gained traction for chart tasks. Examples include UniChart [29], which adapts Donut [17] for chart comprehension, and Pix2Struct [18], which encodes visual features of chart screenshots for direct QA. State-of-the-art systems now harness instruction-tuned multimodal large language models like LLaVA [23], PaliGemma [2, 32], and ChartLlama [9] to process complex queries. Additionally, specialized techniques address numerical reasoning by coupling code execution and large language models (Program-of-Thought), as demonstrated by TinyChart [36] and ChartGemma [30]. These approaches significantly boost answer accuracy but typically provide limited transparency about which regions of the chart were utilized for reasoning. Integrating explicit bounding-box-based visual grounding remains a gap, motivating our work to unify high-level chart QA performance with localized, interpretable evidence.
@@ 3 RefChartQA Benchmark
To enable visual grounding for chart images, we introduce RefChartQA, a benchmark that extends ChartQA [28] and a subset of TinyChart’s ChartQA-PoT [36] with explicit grounding annotations. Unlike prior datasets that focus solely on answer accuracy, RefChartQA ensures models justify their predictions by linking answers to relevant chart elements. Grounding annotations are generated through a multi-stage process, including a heuristic-based extraction, a Program-of-Thought (PoT)-based method, and a GPT-based method. The whole process is summarized in Fig. 3. Next, we discuss how we formed the annotations and present the statistical analysis of our benchmark.
Figure 3: Overview of the stages in the RefChartQA annotation process. A sample is considered successfully annotated by a stage when all its steps are completed. Orange and red downward arrows indicate points where validation failures may lead to sample rejection at that stage.
@@ 3.1 Annotation Process
We implement visual grounding using a bounding box paradigm, which is particularly effective for chart images due to its versatility across different chart types. However, we do not claim this to be the best or only approach; alternative methods, such as semantic segmentation or keypoint detection, could provide complementary advantages by capturing finer details or structural relationships within charts. These methods, however, often come with higher computational costs. For instance, autoregressive segmentation typically requires significantly more processing power and training time. While the keypoint paradigm is a more native approach, it is less suited for many common visualizations where answers correspond to regions rather than single points. Future work could explore multi-grounding paradigms that adapt based on the image and query context. In identifying grounding, we prioritize relevant data values, as they often implicitly indicate the associated structural information. A data element is included in the grounding if it is directly referenced in the answer or is essential for computing it. For min/max questions, we include only the identified element, while for operations like sum or average, all contributing elements are included. When the answer depends on structural aspects rather than specific data values, we annotate the relevant structural element.
We choose ChartQA [28] as the foundation for our dataset due to its diverse question types, covering both arithmetic and non-arithmetic reasoning over chart images. Arithmetic questions often require multi-element operations, such as calculating averages or making comparisons, while non-arithmetic questions typically refer to a single chart element, though some may involve multiple elements for contextual understanding. ChartQA provides comprehensive structured metadata for each chart, including bounding boxes and textual content for both descriptive elements (such as axis labels and legend entries) and data visualization elements within the chart. Data elements are categorized and include color information. Additionally, we incorporate a subset of the ChartQA-PoT dataset [36], as it provides additional arithmetic questions based on ChartQA charts and structured reasoning in the form of Python code. Its use of ChartQA charts allows the reuse of ChartQA’s metadata without modifying the annotation process. However, both datasets lack explicit connections between visual elements and the questions, as they do not include direct grounding annotations. To bridge this gap, we develop a structured annotation process that links questions to their supporting visual evidence as summarized in Fig. 3.
The annotation pipeline is designed to match questions with relevant evidence elements from the metadata. As a result, our dataset may carry over errors from the underlying source datasets, as well as introduce new mistakes during the grounding annotation process. To minimize errors, we incorporate validation in the annotation process.
@@
Single-Element Extraction. The first stage of the annotation pipeline targets single-element non-arithmetic questions, where the answer corresponds directly to a single data element in the chart. Since these questions do not involve computational reasoning or the interaction of multiple elements, a heuristic-based approach is used to match answers with their corresponding visual elements. For example, the question "What is the number of Macy’s stores worldwide in the year 2018?" describes a data retrieval operation which can be answered using structural information only. To identify the relevant evidence element, the question and answer text are analyzed to extract x-axis labels, categories, and data values (from the answer only). String- and color matching, based on metadata, are used to identify the referenced elements. The extracted attributes are used to constrain the set of data element candidates. If a single match remains, it is selected as the grounding element. Otherwise, the annotation fails and the question progresses to the next stage.
@@
PoT Extraction. The second stage introduces PoT-based reasoning to handle more complex question types, particularly those requiring arithmetic operations or multi-element interactions. Unlike the first stage, which relies on direct string-matching, this stage leverages TinyChart’s PoT [36] to reconstruct the logical reasoning path necessary to derive the grounding. For example, consider the following PoT code:
{listing}
@@
The PoT process above starts by extracting the relevant metadata needed to answer the question—specifically, the list of years X and the number of people doing freelance work each year Y. This is followed by a sequence of reasoning operations, culminating in the final computed answer, which is stored in the variable Answer. To ensure consistency with ChartQA annotations, we first align the extracted list values with the chart’s metadata. This step establishes a mapping between approximated values and metadata values while also verifying the accuracy of the extracted elements. To further minimize annotation errors, we enforce that the PoT prediction must exactly match the ground truth answer. After validation, we extract the relevant reasoning elements using a template-based approach. The corresponding metadata elements serve as grounding annotations. If validation fails or the reasoning complexity exceeds the capabilities of the template-based approach, the question advances to the next stage.
@@ 5.1 Model Selection
To evaluate the impact of different pretrained models, we selected several SOTA MLLMs across chart analysis, document understanding, and visual grounding. For chart-related models, we chose UniChart [29], ChartGemma [30], and TinyChart [36]. UniChart, a lightweight Donut-based [17] model with 260M parameters, delivers competitive performance despite its smaller size. In contrast, ChartGemma and TinyChart ($\sim$3B parameters) incorporate advanced reasoning capabilities. TinyChart supports two inference modes: direct inference for end-to-end predictions and PoT-based reasoning, which generates executable Python code. ChartGemma, however, exclusively relies on PoT-based predictions. All three models have been instruction-tuned on chart-related tasks like summarization, chart-to-table conversion, and chart VQA. For visual grounding, we selected Qwen2.5-VL [1] and Qwen-VL-Chat [33], one of the few VG models fine-tuned on ChartQA. Alongside image captioning, VQA, and OCR, it has been trained on grounded image captioning and object referral, making it a strong candidate for grounding-based evaluations. For document understanding, we included DocOwl 1.5 [11], which, like chart-focused models, has been trained on document parsing, table extraction, chart parsing, and VQA, while also incorporating text localization.
@@
5.2 Experiment Setup
We conducted our experiments on four Nvidia A40 GPUs with 48 GB of VRAM. All models were fine-tuned by parameter-efficient fine-tuning using LoRA adapters [12] for five epochs with a global batch size of 16. For UniChart, we used full-parameter instead of LoRA tuning due to its comparatively small parameter count. For the full configuration details, please refer to Table 3. At this point, we treat the bounding box predictions as text and use the next-token prediction objective, with a softmax cross-entropy loss. For inference we relied on direct inference mode and the following input prompt (with slight variances depending on how the model processes images):
“<image>{question} Append grounding bounding boxes.”
For UniChart, a special ground-token task was introduced:
“<grounding_chartqa>{question} <s_answer>”
Table 3: Training configurations for all models.
Model Approach Epochs Batch LR Optimizer Scheduler $\mathbf{r}$ $\bm{\alpha}$ Dropout
UniChart Full param. 5 16 5.00E-05 Adam cosine
ChartGemma LoRA 5 16 1.00E-04 AdamW 8 8 0
TinyChart-PoT LoRA 5 16 1.00E-04 AdamW cosine 64 16 0.05
TinyChart LoRA 5 16 1.00E-04 AdamW cosine 64 16 0.05
QwenVL-Chat LoRA 5 16 1.00E-05 AdamW cosine 64 16 0.05
QwenVL-2.5 LoRA 1 16 2.00E-04 AdamW cosine 16 16 0.05
DocOwl1.5 LoRA 5 16 2.00E-05 AdamW cosine 128 256 0.05
5.3 Evaluation Metrics
@@ 6 Conclusion
In this paper, we introduced RefChartQA, the first benchmark that integrates chart question answering with visual answer grounding to enhance interpretability and response reliability in chart-based reasoning. Through comprehensive experiments on six state-of-the-art vision-language models across different categories, we demonstrate that incorporating grounding improves response accuracy, while also identifying key factors influencing text-spatial alignment. Despite these advancements, challenges remain in handling complex multi-element reasoning and numerical inference. We hope this work will drive further research in chart understanding and VAG, fostering more robust, interpretable, and spatially aware document understanding models.